OpenStudy (anonymous):

7x + 9y = 119 2x + y = 23 substitute

OpenStudy (anonymous):

multiply second equation by 9 and subtract it from first to find value of x, then use that value of x in either equations to find value of y

OpenStudy (anonymous):

isn that eliminiation???

jimthompson5910 (jim_thompson5910):

Step 1) Solve the second equation for y 2x + y = 23 y = 23 -2x ----------------------------------------- Step 2) Plug this into the other equation 7x + 9y = 119 7x + 9(23 - 2x) = 119 7x + 9*23 - 9*2x = 119 7x + 207 - 18x = 119 ...Keep going to solve for x... ----------------------------------------- Step 3) Plug the solution you found in step 2 into y = 23 - 2x and simplify to find y

OpenStudy (anonymous):

@ravergurl , indeed it is elimination, are you supposed to use a particular method ? If yes, please sub y=23-2x in first equation

OpenStudy (anonymous):

ya its substitution... use substitution to solve..(:

OpenStudy (anonymous):

is x=3.52?? doesnt sound right

OpenStudy (anonymous):

x should be 8